# UniCUE: Unified Recognition and Generation Framework for Chinese Cued Speech Video-to-Speech Generation

### v1

@@ Abstract.
Cued Speech (CS) enhances lipreading through hand coding, providing precise speech perception support for the hearing-impaired. CS Video-to-Speech generation (CSV2S) task aims to convert the CS visual expressions (CS videos) of hearing-impaired individuals into comprehensible speech signals. Direct generation of speech from CS video (called single CSV2S) yields poor performance due to insufficient CS data. Current research mostly focuses on CS Recognition (CSR), which convert video content into linguistic text. Based on this, one straightforward way of CSV2S is to combine CSR with a Text-to-Speech system. This combined architecture relies on text as an intermediate medium for stepwise cross-modal alignment, which may lead to error propagation and temporal misalignment between speech and video dynamics. To address these challenges, we propose a novel approach that directly generates speech from CS videos without relying on intermediate text. Building upon this, we propose UniCUE, the first unified framework for CSV2S, whose core innovation lies in the integration of the CSR task that provides fine-grained visual-semantic information to facilitate speech generation from CS videos. More precisely, (1) a novel fine-grained semantic alignment pool to ensure precise mapping between visual features and speech contents; (2) a VisioPhonetic adapter to bridge cross-task representations, ensuring seamless compatibility between two distinct tasks (i.e., CSV2S and CSR); (3) a pose-aware visual processor is introduced to enhance fine-grained spatiotemporal correlations between lip and hand movements in CS video. Experiments on our new established Chinese CS dataset (14 cuers111Cuer represents the person who perform CS.: 8 hearing-impaired and 6 normal-hearing) show that our UniCUE significantly reduces Word Error Rate by 78.3% and improves lip-speech synchronization by 32% compared to the single CSV2S.
@@ 1. Introduction
Cued Speech (CS) is an advanced visual phonetic encoding system that utilizes specific hand shapes and positions to enhance lip reading, providing an accurate visual representation of all phonemes in spoken language (Cornett, 1967; Liu and Feng, 2019; Leybaert and LaSasso, 2010). CS maintains a high level of consistency with spoken language in terms of phonemes and speech patterns, enabling hearing-impaired individuals to better integrate into speech-dominant social and educational environments (Cornett, 1967; Leybaert and LaSasso, 2010; Leybaert et al., 2010). In Mandarin Chinese, CS employs 8 hand shapes and 5 positions to encode consonants and vowels, addressing challenges such as the phonemes with similar lip shapes (Liu and Feng, 2019), as illustrated in Figure 1(a).
@@
CS Video-to-Speech generation (CSV2S) task aims to convert the visual expressions (CS videos) of hearing-impaired individuals into comprehensible speech signals. However, directly constructing an end-to-end CSV2S model faces the following challenges: First, the complex semantic correlation between CS video and semantic content requires precise mapping from visual cues (lip reading and hand coding) to speech, while limited CS dataset scale further constrains model capacity. Second, fine-grained visual spatiotemporal modeling is crucial to resolve the natural asynchrony where hand gestures initiate earlier than corresponding lip movements i.e., hand-preceding phenomenon (Liu et al., 2020). Third, personalized adaptation must address dual variations: (1) inter-cuer differences in CS expression, and (2) acoustic diversity in cuer identity characteristics. To the best of our knowledge, no prior work has addressed CSV2S task.
Existing research primarily focuses on CS Recognition (CSR) that converts CS videos into phoneme-level text(Liu et al., 2020, 2024, 2018, 2019), neglecting the critical need for natural speech generation. This gap severely limits real-time communication between hearing-impaired and normal-hearing individuals, particularly in educational and social contexts. For example, in social gatherings, normal-hearing individuals may need to quickly and accurately understand and respond to questions or information shared by their hearing-impaired friends. However, existing CSR methods can only provide text output, which is insufficient to support natural and smooth interactions. Moreover, lip-reading video-to-speech synthesis methods like LipVoicer (Yemini et al., 2023) leverage lip motions alone, and they fail to model the complementary hand coding in CS. Thus, the primary motivation of this study is to build the first CSV2S system that directly decoding Chinese CS video into speech, as shown in Figure 1(b).
Based on the existing CSR methods, a straightforward approach to build CSV2S is to combine CSR with a Text-to-Speech (TTS) model, as shown in Figure 2(a). However, this combined architecture has fundamental drawbacks. Firstly, the intermediate CS text serve as a medium between visual and speech, inheriting recognition errors from the CSR module, leading to inaccurate speech content. Secondly, the CSR stage only extracts semantic text but loses fine-grained temporal information from video, resulting in generated speech lacking temporal alignment.
@@
Recent advances in multimodal generation have demonstrated that integrating semantic reasoning capabilities from vision-language models (VLMs) can significantly enhance complex generation tasks, such as text-guided image synthesis with interleaved control (Mi et al., 2025; Chen et al., 2025a). Inspired by this paradigm, we believe that the fine-grained visual understanding inherent in CSR can serve as a semantic compensator for CSV2S generation. As illustrated in Figure 2(b), our framework introduces a shared visual processor that bridges CSR and CSV2S tasks. This processor operates as a two-way translator: during CSR, it extracts linguistic semantics from spatiotemporal lip-hand patterns; for CSV2S, it maps these semantics to guide speech synthesis while preserving cuer identity. The key innovation lies in explicitly modeling the semantic compensation flow, where CSR-derived phoneme-level constraints reduce potential ambiguities in word-level speech generation, overcoming the challenges of fine-grained semantic correlation modeling.
@@
To address the above challenges, building upon the semantic compensation paradigm, in this work, we propose UniCUE, the first unified framework that bridges CSR and CSV2S tasks through three specific components: Firstly, Unlike prior CSR methods (Liu et al., 2024; Liu and Liu, 2023) that process lip and hand modalities independently, relying solely on raw video embeddings, UniCUE is equipped with a pose-aware visual processor that simultaneously processes both modalities. This allows UniCUE to generate a mixed visual representation by combining video embeddings and pose embeddings. This mixed setup not only achieves fine-grained spatiotemporal modeling for hand-preceding phenomenon but also enhances robustness to cuer-specific CS expression. Secondly, to enhance the semantic alignment between visual cues and spoken content, we introduce a semantic alignment pool. This mechanism aligns the video latent space and the pose latent space with linguistic semantics via contrastive learning, enabling the model to capture complex semantic correlations and improve the accuracy of the generated speech. Thirdly, we incorporate the visual processor from CSR into the diffusion-based CSV2S model, using an VisioPhonetic Adapter (VPA) to translate fine-grained representation into diffusion-compatible codes. Based on the identity correlation between speech and face portrait (Rong and Liu, 2024), this design enables our diffusion decoder to not only effectively integrate the semantic information conveyed by the visual cues but also capture the unique characteristics of the cuer identity in the visual frames, thereby enhancing accurate cuer-specific speech generation.
@@
To evaluate UniCUE on hearing-impaired individuals, we collect CS data from 8 hearing-impaired cuers. Additionally, we extend the Mandarin Chinese CS dataset (Lei et al., 2024b) by incorporating data from 8 hearing-impaired and 6 normal-hearing cuers (individuals who perform CS), resulting in a total of 11,282 CS videos. The results on this dataset demonstrate that our method not only achieves accurate linguistic understanding but also preserves the cuer’s characteristics.
@@
We propose the first CSV2S generation framework by constructing a new unified multimodal system that integrating CSR capabilities to enhance CSV2S.
@@
We establish a new Mandarin Chinese CS dataset comprising 8 hearing-impaired and 6 normal-hearing cuers. Experimental results demonstrate that our UniCUE framework outperforms the single CSV2S model in terms of speech consistency, accuracy, and time efficiency.
@@ 2.1. Video-to-Speech Generation
V2S involves generating natural-sounding speech synchronized with a silent video of a person talking. A key challenge in this area is performing complex visual-speech modeling under the limited data. To address this challenge, Uni-Dubbing (Lei et al., 2024a) incorporates modality alignment during the pre-training phase on multimodal datasets and then fine-tunes using both multimodal and single-modality audio data. Similarly, Kefalas et al. (Kefalas et al., 2024) leverage abundant audio-only datasets to pre-train the speech encoder and decoder, subsequently fine-tuning with paired speech-video data. Instead of relying on external data, some works (Kim et al., 2023; Yemini et al., 2023; Gupta et al., 2024) have found that incorporating text transcripts plays a significant role in enhancing V2S generation. For example, Kim et al. (Kim et al., 2023) propose guiding the model using multimodal supervision, i.e., text and speech, to complement insufficient word representations via multi-task learning. In contrast to directly using text transcripts, we innovatively utilize the shared semantic correlations between visual-text and visual-audio in CSR and CSV2S.
@@ 2.2. Cued Speech Recognition
CS is an augmented lip-reading system enhanced by hand coding, offering significant support to the hearing-impaired community. While CSR task focuses on converting CS videos into linguistic texts, leveraging both lips and hands as distinct modalities that provide complementary information (Liu and Feng, 2019; Papadimitriou and Potamianos, 2021). Existing CSR methods often process lip and hand features independently, followed by feature fusion for content recognition (Liu et al., 2020, 2024; Liu and Liu, 2023; Zhang et al., 2023). Due to the asynchronous nature of lip and hand movements, achieving effective fusion remains a challenging problem. Liu et al. (Liu et al., 2020) introduced a re-synchronization procedure to align hand features with lip features, addressing the temporal mismatch. While a transformer-based cross-modal mutual learning framework has been developed to facilitate multimodal interaction (Liu and Liu, 2023; Liu et al., 2024). Zhang et al. (Zhang et al., 2023) further explored privacy risks in the training process by applying federated learning. In contrast, our approach takes whole video frames as input that models lip and hand features simultaneously, avoiding the complexities of feature fusion. Moreover, we propose a pose-aware visual processor to enhance feature representation between the two modalities, thereby improving recognition accuracy.
@@ 2.3. Unified Understanding and Generation
Recent efforts in unifying understanding and generation tasks can be broadly categorized into two paradigms. The first paradigm integrates visual-language understanding with external generative models (e.g., diffusion models) to achieve multimodal generation (Wu et al., 2023; Dong et al., 2023; Jin et al., 2023; Li et al., 2024; Sun et al., 2024; Ge et al., 2024b, a). For instance, works like (Jin et al., 2023; Li et al., 2024) leverage large language models (LLMs) for semantic comprehension while relying on diffusion models (Rombach et al., 2022; Podell et al., 2023) for high-quality image synthesis. The second paradigm focuses on optimizing a LLMs-based foundation models through next-token prediction objectives for both visual understanding and generation (Yu et al., 2023; Sun et al., 2023; Zhou et al., 2024; Wu et al., 2024b; Fang et al., 2024; Wu et al., 2024a; Chen et al., 2025b). For instance, Transfusion (Zhou et al., 2024) unifies the processes of image understanding and image generation within a transformer. This approach not only inherits the strong comprehension ability of language models but also preserves the detailed information of images, achieving better controllable generation results in text-to-image. Despite their success, these frameworks predominantly address visual-text scenarios, leveraging understanding for speech generation is unexplored. In contrast to prior works, we present the first unified framework that bridges visual understanding and speech generation.
@@ 3.1. UniCUE
To achieve accurate CSV2S generation, the proposed framework need simultaneously resolve two critical challenges: (1) semantic understanding of the semantic correlations between visual cues and speech content, and (2) speech synthesis that preserves cuer-specific characteristics and temporal alignment. Inspired by the auxiliary benefits of unified understanding and generation for multimodal controllable image synthesis (Wu et al., 2023; Dong et al., 2023), we design a unified architecture that integrates CSR and CSV2S, enabling CSV2S with understanding capability improvement through shared visual feature representations. As illustrated in Figure 3, the framework operates via two pathways.
CSR: Fine-grained Visual Cues Understanding. Given a CS video and its corresponding pose (extracted via OpenPose (Cao et al., 2019)), a pose-aware visual processor extracts the embedding of the visual mixture . The visual embedding is first transcribed into linguistic text via a transformer-based text decoder to model a fine-grained spatiotemporal visual representation space. The process can be denoted as: , where is the predicted text tokens. The learning objective of the CSR path is to predict text tokens as closely as possible to the ground truth text tokens :
where and are the -th token of the predicted tokens and ground truth, respectively, and is the total number of tokens. By transcribing cross-cuer visual cues into linguistic text, CSR path establishes phoneme-level semantic grounding for the visual embedding , which equips with cuer-adaptive, fine-grained semantic understanding capabilities.
CSV2S: Cuer-specific Speech Synthesis. Based on the correlation between face portrait and speech (Rong and Liu, 2024), a latent diffusion model (LDM) (Rombach et al., 2022) is trained to synthesize cuer-specific speech conditioned on the learned visual embedding . Specifically, a pretrained VAE encoder compresses ground-truth mel-spectrograms into latent codes , which are then corrupted into by sequentially injecting Gaussian noise over time steps:
where represents the noise level at the time step. The noised is then input into the LDM (Rombach et al., 2022) to learn the distribution of latent speech codes under the visual condition . The objective of the CSV2S generation path can be formulated as:
where represents the LDM. By training LDM, visual information is interacted with latent speech codes, enabling the generated speech to accurately match the dynamic changes in the visual content while preserving the naturalness of personalized speech.
UniCUE: Unified Understanding and Generation. The CSR path learns fine-grained visual embeddings by performing linguistic text recognition, which are then used as conditional inputs for the CSV2S path, enabling CSV2S to leverage the fine-grained visual understanding for accurate speech generation. This shared mechanism in our unified framework not only reduces information loss but also avoids error propagation that might arise from intermediate text conversions. Therefore, CSV2S can capture cue-specific acoustic features while maintaining linguistic accuracy and temporal consistency, generating cuer-specific speech outputs.

### v3

@@ Abstract.
Cued Speech (CS) enhances lipreading via hand coding, offering visual phonemic cues that support precise speech perception for the hearing-impaired. The task of CS Video-to-Speech generation (CSV2S) aims to convert CS videos into intelligible speech signals. Most existing research focuses on CS Recognition (CSR), which transcribes video content into text. Consequently, a common solution for CSV2S is to integrate CSR with a text-to-speech (TTS) system. However, this pipeline relies on text as an intermediate medium, which may lead to error propagation and temporal misalignment between speech and CS video dynamics. In contrast, directly generating audio speech from CS video (direct CSV2S) often suffer from the inherent multimodal complexity and the limited availability of CS data. To address these challenges, we propose UniCUE, the first unified framework for CSV2S that directly generates speech from CS videos without relying on intermediate text. The core innovation of UniCUE lies in integrating a understanding task (CSR) that provides fine-grained CS visual-semantic cues to to guide the speech generation. Specifically, UniCUE incorporates a pose-aware visual processor, a semantic alignment pool that enables precise visual–semantic mapping, and a VisioPhonetic adapter to bridge the understanding and generation tasks within a unified architecture. To support this framework, we construct UniCUE-HI, a large-scale Mandarin CS dataset containing 11,282 videos from 14 cuers, including both hearing-impaired and normal-hearing individuals. Extensive experiments conducted on this dataset demonstrate that UniCUE achieves state-of-the-art (SOTA) performance across multiple evaluation metrics.
@@ 1. Introduction
Cued Speech (CS) is an visual phonetic encoding system that utilizes specific hand shapes and positions to enhance lip reading, providing an accurate visual representation of all phonemes in spoken language (Cornett, 1967; Liu and Feng, 2019; Leybaert and LaSasso, 2010). CS maintains a high level of consistency with spoken language in terms of phonemes and speech patterns, enabling hearing-impaired individuals to better integrate into speech-dominant social and educational environments (Cornett, 1967; Leybaert and LaSasso, 2010; Leybaert et al., 2010). In Mandarin Chinese, CS employs 8 hand shapes and 5 positions to encode consonants and vowels (as illustrated in Figure 1(a)), addressing challenges such as the phonemes with similar lip shapes (Liu and Feng, 2019).
@@
CS Video-to-Speech generation (CSV2S) task aims to convert CS videos of into comprehensible speech signals. However, directly constructing an end-to-end CSV2S model faces several challenges. Firstly, this task involves complex multimodal semantic correlations, requiring precise mapping from visual cues (lip movements and hand coding) to acoustic speech, while the limited scale of existing CS datasets further constrains model capacity. Secondly, fine-grained spatiotemporal modeling of visual information is essential to resolve the intrinsic asynchrony, i.e., the hand-preceding phenomenon, where hand cues precede corresponding lip movements (Liu et al., 2020a). To the best of our knowledge, the CSV2S task has not been explicitly studied in prior literature.
Existing research primarily focuses on CS Recognition (CSR) that converts CS videos into phoneme-level text (Liu et al., 2020a, 2024, 2018, 2019), neglecting the critical need for natural speech generation. This limitation significantly impairs real-time communication between hearing-impaired and normal-hearing individuals, especially in educational and social scenarios. For instance, in group conversations, normal-hearing participants must quickly comprehend and respond to questions posed by their hearing-impaired peers. Textual output from CSR systems is often insufficient for such natural and smooth interactions. Additionally, recent lipreading-based video-to-speech models such as LipVoicer (Yemini et al., 2024) rely solely on lip movements, failing to capture the complementary hand-coded information in CS that conveys critical phonemic distinctions. These shortcomings underscore the need for a more comprehensive approach. Motivated by this, we aim to develop the first Chinese CSV2S system that directly decodes CS videos into intelligible speech, as illustrated in Figure 1(b).
A straightforward solution, shown in Figure 2(a), is to combine a CSR model with a Text-to-Speech (TTS) system. However, this combined pipeline suffers from two key drawbacks. Firstly, the intermediate textual representation introduces error propagation, as misrecognitions in the CSR stage lead to incorrect speech output. Secondly, the textual intermediate discards fine-grained spatiotemporal cues in the CS video, resulting in synthesized speech that lacks temporal alignment with the visual input.
@@
To overcome these challenges, we draw inspiration from recent advances in multimodal learning, where semantic reasoning from vision-language models (VLMs) has shown strong promise in tasks like text-guided image synthesis with interleaved control (Mi et al., 2025; Chen et al., 2025a). We hypothesize that the multimodal visual understanding inherent in CSR can serve as a semantic bridge to support more accurate and controllable speech generation in CSV2S. As depicted in Figure 2(b), we introduce a unified framework that leverages a shared visual processor to bridge CSR (understanding task) and CSV2S (generation task). This processor serves as a two-way translator: during CSR, it extracts linguistic semantics from fine-grained lip-hand motion patterns; in CSV2S, it utilizes these semantics to guide speech generation. The core innovation of our framework lies in modeling a semantic compensation flow, where phoneme-level supervision from CSR reduces ambiguity in speech synthesis, enabling more faithful and coherent voice generation under complex multimodal conditions.
@@
Building upon this semantic compensation paradigm, in this work, we propose UniCUE, the first unified framework that bridges CSR and CSV2S tasks through three specific components: Firstly, unlike prior CSR methods (Liu et al., 2024; Liu and Liu, 2023) that process lip and hand modalities independently and rely on raw video embeddings, UniCUE employs a pose-aware processor that fuses video and pose streams into a mixed representation. This enables fine-grained spatiotemporal modeling of the hand-preceding phenomenon and improves generalization to cuer-specific expressive styles. Secondly, to enhance the alignment between visual and linguistic semantics, we introduce a semantic alignment pool to map the video and pose latent spaces into a shared textual space using contrastive learning. This facilitates cross-modal correlation modeling and improves semantic consistency in the generated speech. Thirdly, to unify the understanding and generation tasks, we reuse the CSR visual encoder within our diffusion-based CSV2S decoder and introduce a VisioPhonetic Adapter (VPA) that transforms the visual representations into diffusion-compatible codes. This design enables the decoder to effectively incorporate fine-grained semantic information derived from multimodal visual inputs
@@
To evaluate UniCUE on hearing-impaired individuals, we extend the MCCS dataset (Lei et al., 2024b) by adding data from 8 hearing-impaired and 2 normal-hearing cuers111Cuer means the people who perform CS., forming the Unified-HI Corpus with 14 cuers. Experimental results on this dataset demonstrate that UniCUE not only produces accurate and intelligible speech, but also maintains temporal synchronization with the CS video.
@@
We propose the first CSV2S framework by constructing a unified multimodal system that integrates CSR capabilities to enhance speech generation.
@@
We construct a new Mandarin Chinese CS dataset comprising both hearing-impaired and normal-hearing cuers. Experimental results demonstrate that our UniCUE outperforms the state-of-the-art (SOTA) methods in terms of speech accuracy, consistency, and quality.
@@ 2.1. Video-to-Speech Generation
V2S aims to synthesize natural speech aligned with silent talking videos, but is challenged by limited data. Uni-Dubbing (Lei et al., 2024a) addresses this via modality-aligned pre-training on multimodal data and fine-tuning with both multimodal and audio-only inputs. Similarly, Kefalas et al. (Kefalas et al., 2024) pre-train on large audio-only corpora before tuning on paired data. Some studies (Kim et al., 2023; Yemini et al., 2024; Gupta et al., 2024) incorporate transcripts to enhance generation. Kim et al. (Kim et al., 2023) use text-speech supervision to improve word-level representation via multi-task learning. Existing V2S methods primarily focus on lip reading. However, CS conveys phonemic information through both lip and hand movements. Ignoring hand cues results in incomplete visual representations and degraded speech synthesis quality, limiting the applicability of these methods to CS. Notably, no prior work has addressed the CSV2S task.
@@ 2.2. Cued Speech Recognition
CS augments lip reading with hand coding to support the hearing-impaired. The CSR task aims to transcribe CS videos into text by leveraging lips and hands as complementary modalities (Liu and Feng, 2019; Papadimitriou and Potamianos, 2021). Most CSR methods extract lip and hand features separately and fuse them for recognition (Liu et al., 2020a, 2024; Liu and Liu, 2023; Zhang et al., 2023). Due to the asynchronous nature of these modalities, effective fusion remains challenging. Liu et al. (Liu et al., 2020a) proposed re-synchronization to align hand with lip features, while transformer-based mutual learning (Liu and Liu, 2023; Liu et al., 2024) improves multimodal interaction. Zhang et al. (Zhang et al., 2023) addressed privacy concerns via federated learning. In contrast, we directly model lip and hand cues from whole frames, avoiding explicit fusion. A pose-aware visual processor is introduced further to enhance cross-modal representation and improve performance.
@@ 2.3. Unified Understanding and Generation
Recent advances in unifying understanding and generation tasks fall into two main paradigms. The first integrates visual-language understanding with external generative models (e.g., diffusion models) for multimodal generation (Wu et al., 2024a; Dong et al., 2024; Jin et al., 2024; Li et al., 2024; Sun et al., 2024; Ge et al., 2024, 2025). For example, (Jin et al., 2024; Li et al., 2024) utilize large language models (LLMs) for semantic understanding and diffusion models (Rombach et al., 2022; Podell et al., 2023) for high-fidelity image synthesis. The second paradigm trains LLM-based foundation models via next-token prediction for both vision understanding and generation (Yu et al., 2023; Sun et al., 2023; Zhou et al., 2024; Wu et al., 2024b; Fang et al., 2024; Wu et al., 2025; Chen et al., 2025b). Transfusion (Zhou et al., 2024), for instance, unifies image understanding and generation within a single transformer, enabling controllable text-to-image synthesis by preserving visual details. However, existing approaches mainly focus on visual-text settings, leaving visual-to-speech generation underexplored. In this work, we introduce the first unified framework that bridges visual understanding and speech generation.
@@ 3.1. Overview of UniCUE
To achieve accurate CSV2S generation, the proposed method needs to simultaneously address two critical challenges: (1) semantic understanding of the linguistic correlations between visual cues and speech content, and (2) speech synthesis that preserves cuer-specific characteristics and temporal alignment. Inspired by the auxiliary benefits of unified understanding and generation for multi-modal controllable image synthesis (Wu et al., 2024a; Dong et al., 2024), we design a unified architecture that integrates CSR and CSV2S, enabling CSV2S with understanding capability improvement through shared visual feature representations. As illustrated in Figure 3, the framework operates via two pathways.
CSR: Fine-grained Visual Cues Understanding. As the recognition pathway, CSR models fine-grained spatiotemporal visual semantics to transcribe CS videos into linguistic sequences. Given a CS video and its corresponding pose maps (extracted via OpenPose (Cao et al., 2019)), we first utilize a pose-aware visual processor to extract multi-modal embeddings , which capture lip and hand motion cues. And then is fed into a auto-regressive Transformer-based text decoder , which models long-range dependencies and contextual interactions across the sequence to generate the predicted token sequence: , where denotes the predicted token sequence.
Unlike prior approaches relying on Connectionist Temporal Classification (CTC) loss (Graves et al., 2006), which predict each token independently and thus limit the model’s ability to capture cross-token dependencies and coarticulatory effects, our method employs an auto-regressive decoder supervised by cross-entropy loss. This design allows to generate tokens conditioned on previously generated outputs and spatialtemporal visual cues, which is more suited to modeling the asynchronous and dynamic nature of CS.
To further enhance both token-level precision and sequence-level linguistic consistency, we employ a hybrid training objective: a masked language modeling loss supervises selectively masked ground-truth tokens to enhance contextual understanding; a sequence-level cross-entropy loss enforces supervision over the full sequence to promote accurate transcription. The final training objective for CSR is:
where denotes the ground-truth token sequence. This dual-loss strategy enhances token-level accuracy while preserving global sequence semantics, enabling the model to capture subtle visual-linguistic cues and temporal dynamics inherent in CS videos, thus improving recognition performance and supporting speech synthesis.
CSV2S: Cuer-specific Speech Synthesis. To directly synthesize intelligible and personalized speech from CS videos, we formulate speech generation as a conditional denoising process within a latent diffusion model (LDM) (Rombach et al., 2022). Since both lip shapes and hand cues in CS convey phonemic content, the speech generation is conditioned a refined visual embedding , which is derived by transforming the CSR multimodal feature via a VisioPhonetic adapter (VPA). Specifically, a pretrained VAE encoder compresses ground-truth mel-spectrograms into latent codes , which are progressively corrupted with Gaussian noise over steps: where denotes the noise level at timestep . The noisy latent then denoised by the LDM conditioned on . The generation objective is defined as:
where represents the denoising network. By learning this conditional distribution, our model generates temporally aligned speech that reflects the visual expressions of cuers.
UniCUE: Unified Understanding and Generation. The CSR pathway learns fine-grained multi-modal visual embeddings through detailed linguistic recognition. To bridge the architectural gap between the CSR and the diffusion-based speech generator, we introduce a VPA that transforms into a refined representation . These embeddings are subsequently utilized as conditional inputs to the CSV2S pathway, enabling the speech synthesis model to leverage enriched visual understanding for improved generation accuracy. By sharing visual feature representations within this unified framework, our approach effectively reduces information loss and mitigates error propagation that often arises from intermediate text conversions. As a result, CSV2S is capable of generating cue-specific speech that faithfully preserves linguistic fidelity and temporal alignment, producing personalized and intelligible speech outputs tailored to individual cuers.